two train started at the same time from station which were 360 miles apartand travel \ed toward each other, the rate of the fast train exceed the rate of the slow train by 10mph, at the end of 2 hour, they pass each other, find the rate of each train
Answer by checkley75(3666)

2x+2(x+10)=360
2x+2x+20=360
4x=360-20
4x=340
x=340/4
x=85 mph. for the slower train
85+10=95 mph. for the faster train.
Proof:
2*85+2*95=360
170+190=360
360=360